# A Disentangled Representation Learning Framework for Low-altitude Network Coverage Prediction

### v1

@@ V-C2 Network Structure Robustness
To demonstrate the robustness of the proposed algorithm, we visualized the boxplots of different algorithms during ablation experiments, as shown in Fig. 6. The different subplots within the figure visualize varying ranges of the Y-axis. Compared to benchmark 2, the proposed algorithm exhibits a lower mean and smaller fluctuation range. However, benchmark 2 achieves a lower maximum MAE, which may be attributed to the accidentality of outliers. The significantly higher MAE of benchmark 3 underscores the importance of feature compression design. Moreover, we conducted experiments with 20 random splits at low sampling rates of $[1\%,2\%,3\%,4\%,5\%,6\%]$ to simulate early stages of low-altitude network planning with markedly insufficient training data, as well as scenarios where the algorithm is applied across different cities or operators with few matched AAU types between the test and training sets. As illustrated in Fig. 7, it is evident that even at such low training set proportions, the proposed method still achieves acceptable MAE performance (less than $7.5\ \mathrm{dB}$) on the test set for 75% of the cases, significantly outperforming other algorithms. This further highlights the model’s robustness in scenarios with low data availability.
@@ V-C3 Importance of Right Model Knowledge
TABLE IV: Alternative Input for Different Networks in Proposed Method, where Networks 1,2 and 3 for Distance Fading, Frequency Fading, and Antenna Gain Networks, Respectively
@@
Wrong Model 1 $D_{n}$ $\Delta\theta_{n}^{(H)}$ $x^{(S)}_{4}$,$\Delta\theta_{n}^{(V)}$, $\mathbf{x_{4}^{(S)}}$
@@
Based on the model described in Eq. (13), we understand that the sets $\{\Delta\theta_{n}^{(H)},\Delta\theta_{n}^{(V)},\mathbf{x_{4}^{(S)}}\}$, $\{D_{n}\}$, and $\{x^{(S)}_{4}\}$ should be fed into distinct neural networks for learning representations. To investigate the impact of model correctness on learning performance, we considered alternative input combinations for the distance fading, frequency fading, and antenna gain networks, which are denoted as wrong models 1, 2, and 3, respectively, as shown in Table IV. Furthermore, we have illustrated the MAE performance of these models in Fig. 8. It can be observed that the proposed model has an MAE error of less than other wrong models at all BS sampling rates. Compared with the best wrong model 1, the MAE of the proposed model decreases by about $0.2\ \mathrm{dB}$ on average. The observed superior performance of the proposed model highlights the importance of proper model knowledge guidance. Notably, wrong model 1 outperforms wrong models 2 and 3 slightly, which may be attributed to model 1 maintaining a separate network for the $D_{n}$ input, assigning it a greater weight in impact on coverage prediction.
@@
TABLE V: BS Operational Parameters Setting
@@
Figure 10: The prediction coverage visualization results for different operational parameters under different conditions.
Figure 11: SS-RSRP coverage map visualization in Xinyue Lake before operation parameters optimization.
Figure 12: Comparison of predicted and actual coverage at sampling locations in Xinyue Lake before operational parameters optimization (top left: actual SS-RSRP coverage at sampling locations, top right: predicted SS-RSRP coverage at sampling locations, bottom left: absolute difference between actual and predicted SS-RSRP at sampling locations, bottom right: UAV PCI switching during flight).
Figure 13: The distribution of truth-prediction absolute error for all samples in Xinyue Lake before operation parameters optimization.
Figure 14: SS-RSRP coverage map visualization in Xinyue Lake after operation parameters optimization.
Figure 15: Comparison of predicted and actual coverage at sampling locations in Xinyue Lake after operational parameters optimization (top: actual SS-RSRP coverage at sampling locations, bottom: absolute difference between actual and predicted SS-RSRP at sampling locations).
@@ V-D Algorithm Comparison
In addition to the ablation study, the proposed method was compared with several traditional machine learning methods, including KNN[36] (with the number of neighbors set to 50), Lasso [37] (with a regularization strength of 1), Random Forest[38] (with 100 trees), and XGBoost[39] (using mean squared error as the objective). As shown in Fig. 9, the MAE of different algorithms was plotted against varying BS sampling rates. The results indicate that the proposed method consistently outperforms traditional methods across various BS sampling rates, achieving an average MAE improvement of approximately $0.5\ \mathrm{dB}$. This further demonstrates the superiority of the proposed method.
V-E Applications and Visualizations
@@
V-E1 Low-altitude Coverage Prediction for An Individual BS
@@
The effective prediction of low-altitude coverage for individual base stations constitutes a crucial foundation for network coverage. To illustrate the capability of the method in predicting the coverage of a single base station, we have visualized the coverage prediction of the proposed method under different operational parameters as shown in Table V [^4] We only show the different operational parameters. The relative SSB power in Table V is calculated as shown in Eq. (10). In addition, SCENARIO_0 represents a common ground-based wide beam configuration, whereas SCENARIO_21 is configured for aerial beamforming.. Given the extremely limited sample size for individual base stations, it is challenging to quantitatively assess the prediction performance. Therefore, we provide a brief analysis of different coverage prediction results in conjunction with the operational parameters.
@@
From the comparisons in Fig. 10(a) with 10(b) and 10(c), it is evident that the predicted SS-RSRP coverage is better than the predicted 8th-SSB-RSRP coverage. The coverage at a height of 500 meters is slightly weaker than that at 150 meters. From the comparisons in Fig. 10(a) with 10(d), the proposed method effectively captures changes in the horizontal angles of BS parameters and demonstrates better overall low-altitude coverage performance under aerial beam configurations.
V-E2 Multi-BS Coverage Prediction before Operational Parameters Optimization
@@
To identify low-altitude coverage holes, it is often necessary to predict the joint coverage of multiple BSs within a specific area. To achieve this, the proposed method first predicts the coverage of individual BSs. Then, the coverage predictions of multiple BSs are combined and rasterized, with the maximum predicted RSRP in each grid selected as the predicted SS-RSRP coverage for that grid. Using the trained model, we selected an area in the Xinyue Lake region of Jiangxi Province, involving 40 BSs. For each BS, the coverage at a height of 120 meters and a radius of 2000 meters (with a resolution of 10m × 10m) was predicted and combined. To ensure no data leakage, we removed all samples in the training set with longitude in [115.78, 115.794] and latitude in [28.68, 28.69] and trained the proposed model again, as the training set contained some low-altitude UAV measurement samples from the Xinyue Lake region. The visualization results can be shown in Fig. 11.
@@
To evaluate the accuracy of the predictions, additional low-altitude measurements were conducted at a height of 120 meters. The predicted results were sampled at the same locations as the test points and compared with the new measurement results. The visualization results are shown in Fig. 12. It can be observed that the predicted SS-RSRP closely matches the measured RSRP, with the absolute difference between predicted and measured SS-RSRP values being less than $5\ \mathrm{dB}$ for most points. Based on the UAV measurement data of PCI changes for connected BSs, it is evident that larger prediction errors primarily occur at sampling points where BS handovers take place. These sudden handovers introduce prediction errors, which is reasonable. To reduce such errors, future works of coverage modeling and prediction could be incorporated by actual BS handover parameters and cell relationships.
@@
To provide a more precise prediction error, we visualized the distribution of the absolute differences between predicted and measured SS-RSRP values across sampling locations in Fig. 13. The results show over 90% of the sampling have an error of less than $8\ \mathrm{dB}$, and the MAE across all sampling locations is 4.74 dB. This further demonstrates the accuracy and practicality of the proposed method.
V-E3 Multi-BS Coverage Prediction after Operational Parameters Optimization
@@
To enhance the quality of low-altitude coverage in the area, we optimized the operational parameters of 20 BSs in the region following the method described in [26]. Based on the optimized parameters, the coverage at a height of 120 meters was re-predicted and re-measured [^5] We take a new round of low-altitude trajectory collection at 120 meters., as shown in Fig. 14 and 15. Moreover, the coverage prediction error distribution of all samples was visualized in Fig. 16, demonstrating that the proposed algorithm maintained excellent MAE performance. To more effectively demonstrate the algorithm’s effectiveness in adjusting network operational parameters for coverage prediction, we sampled based on the surface coverage in Fig. 11 and calculated the MAE for the adjusted sampling points, as illustrated in Fig. 17. A comparison between Fig. 16 and Fig. 17 reveals that the accuracy of surface coverage prediction based on adjusted operational parameters outperforms the accuracy before adjustment, showing an approximate 18% performance improvement. This indicates the proposed algorithm’s sensitivity to parameter adjustments, allowing for highly accurate coverage evaluation after network optimization.
V-E4 Multi-BS Coverage Prediction in Ganzhou without AAU Type of BSs
@@
When predicting low-altitude coverage across different operators, cities, or equipment manufacturers, the AAU type of the BSs to be predicted is often unknown during model training. To verify the performance of the proposed method in this case, we selected BSs along a low-altitude flight path in Ganzhou, Jiangxi Province, as the subjects for prediction. Given that the BSs in Ganzhou are all manufactured by ZTE, while those in the training data are from Huawei, all the AAU types of the BSs to be predicted are considered as a new category. Consequently, we excluded the AAU type feature $x_{1}^{(S)}$ from the input features and retrained a model using the training set. By inputting the operational parameters of the BSs (a total of 60 BSs) along the flight path, we predicted the low-altitude coverage at 150 meters altitude. Specifically, we predicted the RSRP for each of the 8 beams of every BS, followed by stacking and rasterization, with the maximum predicted RSRP in a grid defined as the grid SS-RSRP. The predicted surface coverage is shown in Fig. 18. Further sampling of the surface coverage, as illustrated in Fig. 19, reveals that the majority of the sampling points along the flight path have a predicted error of less than $5\ \mathrm{dB}$. To quantify the error more accurately, as shown in Fig. 20, we visualized the distribution of the absolute differences between predicted SS-RSRP and real sampled SS-RSRP. The method proposed in this study achieved an MAE of $6.26\ \mathrm{dB}$ and a MAPE of 7.96%, demonstrating the reliability of the method and its transferability across different operator scenarios.
@@
Figure 20: The distribution of truth-prediction error for all samples in Ganzhou after operation parameters optimization.

### v2

@@ V-C2 Network Structure Robustness
To demonstrate the robustness of the proposed algorithm, we visualized the boxplots of different algorithms during ablation experiments, as shown in Fig. 6. The different subplots within the figure visualize varying ranges of the Y-axis. Compared to benchmark 2, the proposed algorithm exhibits a lower mean and smaller fluctuation range. However, benchmark 2 achieves a lower maximum MAE, which may be attributed to the accidentality of outliers. The significantly higher MAE of benchmark 3 underscores the importance of feature compression design. Moreover, we conducted experiments with 20 random splits at low sampling rates of $1\%,2\%,3\%,4\%,5\%,6\%]$ to simulate early stages of low-altitude network planning with markedly insufficient training data, as well as scenarios where the algorithm is applied across different cities or operators with few matched AAU types between the test and training sets. As illustrated in Fig. 7, it is evident that even at such low training set proportions, the proposed method still achieves acceptable MAE performance (less than 7.5dB7.5\ $\mathrm${dB}) on the test set for 75% of the cases, significantly outperforming other algorithms. This further highlights the model’s robustness in scenarios with low data availability.
@@ V-C3 Importance of Right Model Knowledge
TABLE V: Alternative Input for Different Networks in Proposed Method, where Networks 1,2 and 3 for Distance Fading, Frequency Fading, and Antenna Gain Networks, Respectively
@@
Wrong Model 1 $D_{n}$ $\Delta\theta_{n}^{(H)}$ $x^{(S)}_{4}$$\Delta\theta_{n}^{(V)}$, $\mathbf{x_{4}^{(S)}}$
@@
Based on the model described in Eq. (13), we understand that the sets $\{\Delta\theta_{n}^{(H)},\Delta\theta_{n}^{(V)},\mathbf{x_{4}^{(S)}}\}$, $\{D_{n}\}$, and $\{x^{(S)}_{4}\}$ should be fed into distinct neural networks for learning representations. To investigate the impact of model correctness on learning performance, we considered alternative input combinations for the distance fading, frequency fading, and antenna gain networks, which are denoted as wrong models 1, 2, and 3, respectively, as shown in Table V. Furthermore, we have illustrated the MAE performance of these models in Fig. 8. It can be observed that the proposed model has an MAE error of less than other wrong models at all BS sampling rates. Compared with the best wrong model 1, the MAE of the proposed model decreases by about 0.2dB0.2\ $\mathrm${dB} on average. The observed superior performance of the proposed model highlights the importance of proper model knowledge guidance. Notably, wrong model 1 outperforms wrong models 2 and 3 slightly, which may be attributed to model 1 maintaining a separate network for the $D_{n}$ input, assigning it a greater weight in impact on coverage prediction.
@@
V-C4 Algorithm Generalization Capability
Since we do not have infinite unknown data and cannot confirm the distribution of infinite data sets, the theoretical learning/generalization capability cannot be calculated, but we can verify the model’s generalization capability experimentally. In the setting of the training set, validation set, and test set ratio of 6:2:2, 200 models were trained by randomly dividing the data set under 200 random seeds. We visualised the MAE CDF curves of these models in the training set, validation set, and test set.
According to the Fig. 9, the training set of all models is below 3.5 $\mathrm{dB}$, and the median is 2.733 $\mathrm{dB}$, indicating that the model has been basically fully trained. The CDF curves of validation MAE and test MAE are very close, and the median is 5.196 $\mathrm{dB}$ and 5.166 $\mathrm{dB}$, respectively, which are both low, indicating the reliability of the proposed method’s Generalization capability. About 20% of the models’ test MAE is greater than 7 $\mathrm{dB}$, which may be caused by the large gap in data distribution between the training set and the test set. This problem can be improved by appropriately increasing the proportion of the verification set and training multiple models to vote.
Figure 9: CDF curve of MAE metric for 200 trained neural networks across different dataset split seeds.
@@ V-C4 Algorithm Generalization Capability
TABLE VI: Computational Complexity Comparison across Different Algorithms
Metric Proposed KNN Lasso RandomForest XGBoost
Model Size (MB) 4.31 9.27 0.010 172.56 2.66
Parameters $\sim$1.1M N/A* $\sim$2K $\sim$45M $\sim$680K
Inference Time (s)** <<0.5 1.2 0.1 3.8 0.8
Memory Usage (GB) 0.8 2.1 0.05 4.2 1.3
*KNN is instance-based without explicit parameters
**For 2,500 base station predictions on RTX 4090
TABLE VII: BS Operational Parameters Setting
@@
Figure 11: The prediction coverage visualisation results for different operational parameters under different conditions.
Figure 12: Multi-BS coverage prediction framework. Operational parameters from $B$ interested base stations are input to the trained model to generate individual coverage predictions, which are then fused to produce the integrated coverage map.
Figure 13: SS-RSRP coverage map visualisation in Xinyue Lake before operation parameters optimisation.
Figure 14: Comparison of predicted and actual coverage at sampling locations in Xinyue Lake before operational parameters optimisation (top left: actual SS-RSRP coverage at sampling locations, top right: predicted SS-RSRP coverage at sampling locations, bottom left: absolute difference between actual and predicted SS-RSRP at sampling locations, bottom right: UAV PCI switching during flight).
Figure 15: The distribution of truth-prediction absolute error for all samples in Xinyue Lake before operation parameters optimisation.
Figure 16: SS-RSRP coverage map visualisation in Xinyue Lake after operation parameters optimisation.
Figure 17: Comparison of predicted and actual coverage at sampling locations in Xinyue Lake after operational parameters optimisation (top: actual SS-RSRP coverage at sampling locations, bottom: absolute difference between actual and predicted SS-RSRP at sampling locations).
@@ V-D Algorithm Comparison
In addition to the ablation study, the proposed method was compared with several traditional machine learning methods, including KNN[36] (with the number of neighbors set to 50), Lasso [37] (with a regularization strength of 1), Random Forest[38] (with 100 trees), and XGBoost[39] (using mean squared error as the objective). As shown in Fig. 10, the MAE of different algorithms was plotted against varying BS sampling rates. The results indicate that the proposed method consistently outperforms traditional methods across various BS sampling rates, achieving an average MAE improvement of approximately 0.5dB0.5\ $\mathrm${dB}. This further demonstrates the superiority of the proposed method.
The superior performance stems from how the proposed framework addresses fundamental limitations of traditional algorithms in low-altitude scenarios. KNN’s degradation with increased sampling rates occurs because it relies on Euclidean distance in a feature space with imbalanced dimensionality—high-dimensional BS parameters are sparsely sampled while low-dimensional locations vary densely. In low-altitude networks where beam patterns dominate coverage, nearby neighbours in feature space may belong to BSs with drastically different antenna configurations, causing poor predictions. Linear methods like Lasso fail because they cannot capture the non-linear interactions between distance fading, frequency-dependent attenuation, and directional antenna gain described in Eq. (9) and Eq. (13), particularly the angular transformations in Eq. (4)-(6). Tree-based ensembles (Random Forest, XGBoost), while non-linear, process features independently through recursive partitioning without leveraging propagation physics.
In contrast, our disentangled representation network explicitly decomposes predictions into physically meaningful components through dedicated sub-networks for distance fading, frequency fading, and antenna gain, with outputs combined additively per Eq. (13). The expert knowledge-based feature compression (Eq. (4)-(8)) transforms imbalanced features into balanced representations, while the constrained architecture reduces overfitting risk (Propositions 1-2)—critical when individual BS data is sparse. This demonstrates the framework’s superiority in exploiting domain knowledge for improved accuracy and generalisation in low-altitude scenarios.
V-E Computational Complexity Analysis
To assess the practical applicability of our proposed method, we conduct a comprehensive computational complexity analysis from both theoretical and empirical perspectives.
V-E1 Theoretical Complexity Analysis
We first analyze the theoretical computational advantages of our disentangled representation network compared to conventional MLP architectures. Let $V_{\text{net}}$ and $D_{\text{net}}$ denote the hidden layer width and depth, respectively.
For our disentangled representation network, the parameter count in the hidden layer is
$N_{\text{proposed}}=D_{\text{net}}\times\left[\left(\frac{V_{\text{net}}}{3}\right)^{2}+\frac{V_{\text{net}}}{3}\right]\times 3=\frac{V_{\text{net}}^{2}D_{\text{net}}}{3}+V_{\text{net}}D_{\text{net}}.$ (19)
In contrast, a conventional MLP with equivalent representation capacity requires
$N_{\text{MLP}}=V_{\text{net}}^{2}D_{\text{net}}+V_{\text{net}}D_{\text{net}}.$ (20)
This corresponds to a parameter reduction of
$\frac{N_{\text{MLP}}-N_{\text{proposed}}}{N_{\text{MLP}}}=\frac{2V_{\text{net}}^{2}D_{\text{net}}}{3(V_{\text{net}}^{2}D_{\text{net}}+V_{\text{net}}D_{\text{net}})}\approx 66\%.$ (21)
This theoretical analysis demonstrates that by strategically incorporating communication domain knowledge through decoupled spatial processing, our architecture achieves substantial parameter efficiency while maintaining representational capability.
V-E2 Empirical Complexity Comparison
To validate the practical computational advantages, we compare our method with established baseline algorithms across multiple complexity metrics. Table VI presents comprehensive comparisons including model size, parameter count, and inference time.
The empirical analysis reveals that our proposed method achieves an optimal balance between model complexity and computational efficiency. While Lasso demonstrates the smallest model size due to its linear nature, it lacks the representational capacity for complex spatial relationships in LANC prediction. Conversely, RandomForest achieves competitive accuracy but at the cost of significantly larger model size and inference time. Our approach maintains moderate model size (4.31 MB) while delivering sub-second inference for large-scale predictions, making it practically viable for real-time network planning applications.
V-F Applications and Visualizations
@@
V-F1 Low-altitude Coverage Prediction for An Individual BS
@@
The effective prediction of low-altitude coverage for individual base stations constitutes a crucial foundation for network coverage. To illustrate the capability of the method in predicting the coverage of a single base station, we have visualized the coverage prediction of the proposed method under different operational parameters as shown in Table VII [^5] We only show the different operational parameters. The relative SSB power in Table VII is calculated as shown in Eq. (10). In addition, SCENARIO_0 represents a common ground-based wide beam configuration, whereas SCENARIO_21 is configured for aerial beamforming.. Given the extremely limited sample size for individual base stations, it is challenging to quantitatively assess the prediction performance. Therefore, we provide a brief analysis of different coverage prediction results in conjunction with the operational parameters.
@@
From the comparisons in Fig. 11(a) with 11(b) and 11(c), it is evident that the predicted SS-RSRP coverage is better than the predicted 8th-SSB-RSRP coverage. The coverage at a height of 500 meters is slightly weaker than that at 150 meters. From the comparisons in Fig. 11(a) with 11(d), the proposed method effectively captures changes in the horizontal angles of BS parameters and demonstrates better overall low-altitude coverage performance under aerial beam configurations.
V-F2 Multi-BS Coverage Prediction before Operational Parameters Optimization
@@
To identify low-altitude coverage holes, it is often necessary to predict the joint coverage of multiple BSs within a specific area. As illustrated in Fig. 12, the proposed method first predicts the coverage of individual BSs using their respective operational parameters. Then, the coverage predictions of multiple BSs are combined and rasterized, with the maximum predicted RSRP in each grid selected as the predicted SS-RSRP coverage for that grid. Using the trained model, we selected an area in the Xinyue Lake region of Jiangxi Province, involving 40 BSs. For each BS, the coverage at a height of 120 meters and a radius of 2000 meters (with a resolution of 10m × 10m) was predicted and combined. To ensure no data leakage, we removed all samples in the training set with longitude in [115.78, 115.794] and latitude in [28.68, 28.69] and trained the proposed model again, as the training set contained some low-altitude UAV measurement samples from the Xinyue Lake region. The visualization results can be shown in Fig. 13.
@@
To evaluate the accuracy of the predictions, additional low-altitude measurements were conducted at a height of 120 meters. The predicted results were sampled at the same locations as the test points and compared with the new measurement results. The visualization results are shown in Fig. 14. It can be observed that the predicted SS-RSRP closely matches the measured RSRP, with the absolute difference between predicted and measured SS-RSRP values being less than 5dB5\ $\mathrm${dB} for most points. Based on the UAV measurement data of PCI changes for connected BSs, it is evident that larger prediction errors primarily occur at sampling points where BS handovers take place. These sudden handovers introduce prediction errors, which is reasonable. To reduce such errors, future works of coverage modelling and prediction could incorporate actual BS handover parameters and cell relationships.
@@
To provide a more precise prediction error, we visualized the distribution of the absolute differences between predicted and measured SS-RSRP values across sampling locations in Fig. 15. The results show over 90% of the sampling have an error of less than 8dB8\ $\mathrm${dB}, and the MAE across all sampling locations is 4.74 dB. This further demonstrates the accuracy and practicality of the proposed method.
V-F3 Multi-BS Coverage Prediction after Operational Parameters Optimization
@@
To enhance the quality of low-altitude coverage in the area, we optimized the operational parameters of 20 BSs in the region following the method described in [26]. Based on the optimized parameters, the coverage at a height of 120 meters was re-predicted and re-measured [^6] We take a new round of low-altitude trajectory collection at 120 meters., as shown in Fig. 16 and 17. Moreover, the coverage prediction error distribution of all samples was visualized in Fig. 18, demonstrating that the proposed algorithm maintained excellent MAE performance. To more effectively demonstrate the algorithm’s effectiveness in adjusting network operational parameters for coverage prediction, we sampled based on the surface coverage in Fig. 13 and calculated the MAE for the adjusted sampling points, as illustrated in Fig. 19. A comparison between Fig. 18 and Fig. 19 reveals that the accuracy of surface coverage prediction based on adjusted operational parameters outperforms the accuracy before adjustment, showing an approximate 18% performance improvement. This indicates the proposed algorithm’s sensitivity to parameter adjustments, allowing for highly accurate coverage evaluation after network optimization.
V-F4 Multi-BS Coverage Prediction in Ganzhou without AAU Type of BSs
@@
When predicting low-altitude coverage across different operators, cities, or equipment manufacturers, the AAU type of the BSs to be predicted is often unknown during model training. To verify the performance of the proposed method in this case, we selected BSs along a low-altitude flight path in Ganzhou, Jiangxi Province, as the subjects for prediction. Given that the BSs in Ganzhou are all manufactured by ZTE, while those in the training data are from Huawei, all the AAU types of the BSs to be predicted are considered as a new category. Consequently, we excluded the AAU type feature $x_{1}^{(S)}$ from the input features and retrained a model using the training set.
The maintained prediction accuracy without AAU type information can be attributed to two key factors. First, as categorised in Table II, the AAU type primarily determines the availability of antenna channels, which is already explicitly provided as a separate input feature ($x_{2}^{(S)}$, number of channels). The coverage scenario feature ($x_{3}^{(S)}$) further specifies the beam antenna patterns configured for each scenario at the corresponding carrier frequency ($x_{4}^{(S)}$). Therefore, the critical beam characteristics, including beam count, beam width, and beamforming capabilities, are still implicitly captured through the combination of these remaining features without explicit AAU manufacturer-specific information. Second, the disentangled representation framework learns to extract antenna gain patterns through the antenna gain network, which takes $\{\Delta\theta_{n}^{(H)},\Delta\theta_{n}^{(V)},x_{2}^{(S)},x_{3}^{(S)}\}$ as inputs. This network can generalize across different AAU types by learning the functional relationship between relative angles and gain, conditioned on the remaining static beam characteristics. The model essentially learns that BSs with similar channel configurations and coverage scenarios exhibit similar beam patterns, regardless of the specific AAU manufacturer.
@@
Figure 22: The distribution of truth-prediction error for all samples in Ganzhou.
By inputting the operational parameters of the BSs (a total of 60 BSs) along the flight path, we predicted the low-altitude coverage at 150 meters altitude. Specifically, we predicted the RSRP for each of the 8 beams of every BS, followed by stacking and rasterization, with the maximum predicted RSRP in a grid defined as the grid SS-RSRP. The predicted surface coverage is shown in Fig. 20. Further sampling of the surface coverage, as illustrated in Fig. 21, reveals that the majority of the sampling points along the flight path have a predicted error of less than 5dB5\ $\mathrm${dB}. To quantify the error more accurately, as shown in Fig. 22, we visualized the distribution of the absolute differences between predicted SS-RSRP and real sampled SS-RSRP. The method proposed in this study achieved an MAE of 6.26dB6.26\ $\mathrm${dB} and a MAPE of 7.96%. The slight performance degradation compared to the in-domain prediction (MAE of 4.74dB4.74\ $\mathrm${dB} in Fig. 15) is expected due to potential manufacturer-specific differences in beam pattern implementations. Nevertheless, this result demonstrates the reliability of the method and its transferability across different operator scenarios, even when complete BS hardware information is unavailable.